OpenStudy (anonymous):

a rectangle is 6 times as long is at it is wide. the perimeter is 120 what are the dimensions of the rectangle

OpenStudy (anonymous):

If x is the width of the rectangle and y the lenght, then \[y=6x\] and \[2x+2y=120\]\[x+y=60\] Solving the system we get \[6x +x =60\]\[x=60/7\]So \[y=6*60/7=360/7\]